OpenStudy (anonymous):

my 15 yr old grandson is having problems simplifying expressions. example: -(-8-6t). He soesn't have a clue!

jigglypuff314 (jigglypuff314):

distributive property: a( b + c ) = a*b + a*c -1(-8 -6t) = (-1)(-8) + (-1)(-6t) = 8 + 6t

OpenStudy (anonymous):

so if that is the casse for that problem, would -(6+d) = (-1)(6)+(-1)(d)

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

thank you so very much

jigglypuff314 (jigglypuff314):

your welcome :) glad I could help :D